# biblio.bib

### v3

@@ 1 Introduction
Directed Acyclic Graphs (DAGs for short) are directed graphs in which there is no directed path (sequence of incident edges) from a vertex to itself. They are an omnipresent data structure in various areas of computer science and mathematics. In concurrency theory for instance, scheduling problems usually define a partial order on a number of tasks, which is naturally encoded as DAG via its Hasse diagram [CMPTVW2010, CEH2019]: each task corresponds to a vertex in the graph and task dependencies are materialised by directed edges. Scheduling then corresponds to finding a good topological order on this graph. Natural question such as counting or sampling the schedulings of a program are studied in this context for the purpose of random testing [GS2005]. DAGs also appear as the result of the compression of some tree-like structures such as XML documents [BLMN2015]. In functional programming in particular, this happens at the memory layout level of persistent tree-like values, where the term “hash-consing” has been coined to refer to this compression [Goto1974]. Computer algebra systems also make use of this idea to store their symbolic expression [Ershov1958]. These tree use cases leverage the fact that actual memory gains can be obtained by compacting trees, which has been quantified in [FSS1990] and has motivated the study of compacted trees in the recent years [EFW2021, GW2024]. Finally, complex histories, such as those used in version control systems (see Git for instance [git2018, p. 17]) or genealogy “trees” are DAGs as well.
@@ 1.1 Background on DAGs
The problem of counting DAGs has been solved in early 70’s by Robinson and Stanley using different approaches. In [Robinson1970], Robinson exhibits a recursive decompositions of labelled DAGs leading to a recurrence satisfied by the numbers $A_{n,k}$ of DAGs with $n$ vertices including $k$ sources (vertices without any incoming edge). He later reformulates those results in terms of a new kind of generating functions, now called graphical generating functions in [Robinson1973], and also obtains the asymptotic number of size $n$ DAGs. Around the same time, Stanley also used a generating function approach in [Stanley1973] obtained the same results by deriving identities of the chromatic polynomial. Robinson also solves the unlabelled case starting from the same ideas but using Burnside’s lemma and cycle index sums to account for the symmetries of these objects. He provides a first solution in [Robinson1970] and makes it more computationally tractable in [Robinson1977]. In the 90’s, Gessel generalised those results, also using the graphical generating function framework in [Gessel1995, Gessel1996] to take into account more parameters and count DAGs by vertices and edges, but also sinks and sources.
@@
From the point of view of uniform random generation, the recursive decomposition exhibited by Robinson in [Robinson1973] is interesting as it is amenable to the recursive method pioneered by Nijenhuis and Wilf in [NW1978]. This yields a polynomial time algorithm for sampling uniform DAGs with $n$ vertices. The analysis of this algorithm has been done in [KM2015] but it had been acknowledged earlier in [MDB2001] although the article proposes an alternative solution. Both [KM2015] and [MDB2001] also offer a Markov chain approach to the random sampling problem and an interesting discussion on the pros and cons of both approaches is given in [KM2015]. Remote from the field of combinatorics, the random generation of DAGs is also an active topic in the area of applied statistics and Bayesian inference. In this context, DAGs encode a relevant structure in a collection of random variables and the problem of interest is to sample DAGs from a particular distribution related to those random variables. To this end, authors resort both to Monte Carlo Markov Chains approaches [KSM2022, KM2017] and methods similar to what is referred to as the recursive method in combinatorics [TVK2020]. An important point in [KM2017] is better performance can be achieved by using a combination of both approaches, in particular by exploiting the combinatorial properties of DAGs. Notable is that sampling from the uniform distribution is tackled as a particular case in [TVK2020] and solved with the asymptotically optimal $O(n^{2})$ complexity at the expense of a $O(n^{3})$ pre-processing step.
@@
Another interesting question is that of controlling the number of edges in those random samplers. Indeed, sampling a uniform DAG with a prescribed number of vertices and edges cannot be achieved using the Markov chain approach as it constrains the chain too much, and the formulas of Gessel are not amenable to this either. In [KM2015, § 7], the authors provide an interesting discussion on which kind of restrictions can be made on DAGs with the Markov chain approach. They address in particular the case of bounding the number of edges and highlight that the Monte Carlo Markov Chain approach fails when the desired number of edges is too low, thus advocating for having precise combinatorial enumerations.
@@ 1.2 Contributions
In the present paper, we propose to study an alternative model of DAGs, which we call Directed Ordered Acyclic Graphs (DOAGs), and which are enriched with additional structure on the edges. More precisely, a DOAG in an unlabelled DAG where (1) set of outgoing edges of each vertex is totally ordered and (2) the sources are totally ordered as well. This local ordering of the outgoing edges allows to capture more precisely the structure of existing mathematical objects. For instance, the compressed formulas and tree-like structures mentioned earlier (see [Ershov1958, Goto1974]) indeed present with an ordering as soon as the underlying tree representation is ordered. This is the case for most trees used in computer science (e.g. red-black trees, B-trees, etc.) and for all formulas involving non-commutative operators. The model we introduce thus allows for a more faithful modelling of a wide range of objects. We present here several results regarding DOAGs, as well as an extension of our method to classical labelled DAGs.
As a first step of our analysis, we describe a recursive decomposition scheme that allows us to study DOAGs using tools from enumerative combinatorics. This allows us to obtain a recurrence formula for counting them, as well as a polynomial-time uniform random sampler, based on the recursive method from [NW1978], giving full control over their number of vertices and edges. Our decomposition is based on a “vertex-by-vertex” approach, that is we remove one vertex at a time and we are able to describe exactly what amount of information is necessary to reconstruct the graph. This differs from the approach of Robinson to study DAGs, where all the sources of a DAG are removed at once instead. Although this is a minor difference, our approach allows us to easily account for the number of edges of the graph, which is why our random sampler is able to target DOAGs with a specific number of edges. In terms of application, this means that we are able to efficiently sample DOAGs of low density. A second by-product of our approach is that it makes straightforward to bound the out-degree of each vertex, thus allowing to sample DOAGs of low degree as well.
In order to show the applicability of our method, we devise a similar decomposition scheme for counting labelled DAGs with any number of vertices, edges, and sources. This allows us to transfer our results on DOAGs in the context of labelled DAGs. Our new recurrence differs from the formula of Gessel [Gessel1996] in that it does not resort to the inclusion-exclusion principle. Our approach allows us to obtain an efficient uniform random sampler of labelled DAGs with a prescribed number of vertices, edges, and sources. Here again, in addition to giving control over the number of edges of the produced objects, our approach can also be adapted to bound the out-degree of their vertices. To our knowledge, this is the first such sampler.
@@
In terms of applications, our random generation algorithms enable to experiment with the properties of the objects they model and with the average complexity of algorithms operating on them. A similar approach is for instance taken in [CP2024] where samplers for a realistic class of Git graphs are developed in order to tackle the average complexity of a new algorithm introduced in [Lecoq2024, CDL2022]. Random testing a also an important application of random sampling, especially as a building block for property-based testing, a now well-established framework pioneered Claessen and Hughes in [CH2000].
This paper extends an earlier article [GPV2021] with new results on the asymptotics of DOAGs, with an optimal uniform random sampler for the case when the number of edges is not prescribed, and covers a larger class of DOAGs and DAGs by drooping a constraint on the number of sinks. For the sake of completeness, the most important results and ideas from [GPV2021] will be recalled in the present paper, but the reader will have to refer the earlier article to get the full proof and algorithmic details.
@@ 1.3 Outline of the paper
In Section 2, we start by introducing the class of Directed Ordered Acyclic Graphs and their recursive enumeration and describe a recursive decomposition scheme allowing to count them. In Section 3, we quickly go over earlier results regarding the random generation of DOAGs with a prescribed number of vertices, edges, and sources. The presentation given in this paper slightly generalises over the algorithm given in [GPV2021] but the ideas and proofs remain unchanged. Section 4 shows that our approach applies to labelled graphs as well and opens the way for further research regarding this class. We show that our method, when applied to labelled DAGs, yields a constructive counting formula for them, that is amenable to efficient uniform random generation with full control on the number of edges. Then, in Section 5, we present a bijection between DOAGs and class of integer matrices. This bijection is the key result of this paper as it allows to understand the structure of DOAGs in detail, and to obtain both asymptotic and algorithm results in the following sections. In Section 6, we present a first asymptotic result: we give an asymptotic equivalent of the number of DOAGs of size $n$ with any number of sources and edges. We also state some simple structural properties of those DOAGs. In light of the matrix encoding and these asymptotic results, we design an optimal uniform random sampler of DOAGs with a given number of vertices (but no constraint on the number of edges), that is described in Section 7.
@@ 2 Definition and recursive decomposition
In this section, we recall a model of directed acyclic graphs called “Directed Ordered Acyclic Graphs” (or DOAGs) that we introduced in [GPV2021]. It is similar to the classical model of unlabelled DAGs but where, in addition, we have a total order on the outgoing edges of each vertex. The presentation we opted for here slightly differs from that of [GPV2021] but essentially defines the same objects, the only difference being that we now allow several sinks for the sake of generality.
@@
  • for all $v\in V$$\prec_{v}$ is a total order over the set of outgoing edges of $v$;
@@
We study this class as a whole, however, some sub-classes are also of special interest, in particular for the purpose of modelling compacted data structures. Tree structures representing real data, such as XML documents for instance [BLMN2015], are rooted trees. When these trees are compacted, the presence of a root translates into a unique source in the resulting DOAG. Similarly, DOAGs with a single sink will arise naturally when compacting trees which bear a single type of leaves. In particular the model of compacted binary trees, which can also be seen as a class of cycle-free binary automata, has been shown have unusual combinatorial properties in [EFW2020, EFW2021] and corresponds to a restriction of our model with only binary nodes (and one sink). For these reasons, we will also discuss how to approach the sub-classes of DOAGs with a single source and/or a single sink in this document.
@@
Figure 1: All DOAGs with respectly $1$ vertex, $2$ vertices, $3$ vertices, and simultaneously $4$ vertices and $3$ edges. All edges are implicitly oriented from top to bottom, the blue labels and arrows represent the sources and out-edges orderings (always from left to right).
@@ 2.1 Recursive decomposition
  • $I$ is any subset of the internal vertices of $D^{\prime}$ (children of $v$ in $D$);
  • $s$ is any integer between 00 and the number of sources of $D^{\prime}$;
@@
$D_{n,m,k}$ =$=\begin{cases}0&\text{when}~{}k\leq 0\\ \sum_{p=0}^{n-k}\sum_{i=0}^{p}D_{n-1,m-p,k-1+p-i}\binom{n-k-p+i}{i}\binom{p}{i}i!&\text{otherwise,}\end{cases}$
where $p=s+i$ corresponds the out-degree of the smallest source. The term $\binom{n-k-p+i}{i}=\binom{n-k-s}{i}$ accounts for the choice of the set $I$ and the term $\binom{p}{i}i{!}$ accounts for the number of injective functions $f:I\to\llbracket 1;p\rrbracket$. The upper bound on $p$ in the sum is justified by the fact that the out-degree of any vertex can be at most the number of non-sources in the graph, that is $(n-k)$.
The decomposition scheme presented here differs from the approach described by Robinson in [Robinson1970] as it operates on only one source at a time. It is also reminiscent of the peeling processes used in map enumeration where maps are decomposed one face at a time, see for instance [KMSW2019]. However, the absence of ordering amongst the incoming edges of each vertex in our setup renders those approaches inapplicable as is.
@@ 2.1.1 Special sub-classes based on out-degree constraints
Since $p=i+s$ is the out-degree of the removed source in the above summation, it is easy to adapt this sequence for counting DOAGs with constraints on the out-degree of the vertices. For instance, DOAGs with only one sink are obtained by ensuring that every vertex has out-degree at least one. In other words, let the summation start at $p=1$. Note that restricting DOAGs to have only one single sink or one single source ensures that they remain connected, however not all connected DOAGs are obtained this way. As another example, DOAGs with out-degree bounded by some constant $d$ are obtained by letting $p$ range from 00 to $\min(n-k,d)$.
@@
$D^{\mathcal{P}}_{n,m,k}$ =$=\begin{cases}0&\text{when}~{}k\leq 0\\ \sum_{p\in\mathcal{P}}\sum_{i=0}^{p}D^{\mathcal{P}}_{n-1,m-p,k-1+p-i}\binom{n-k-p+i}{i}\binom{p}{i}i!&\text{otherwise,}\end{cases}$
@@ 3 Earlier results on counting and recursive sampling
In this section we summarise our earlier results on the counting and random sampling problem for DOAGs when all three parameters (number of vertices, edges and sources) are fixed. The theorems are stated in a slightly more general setting here than in [GPV2021] so as to capture all variants of the model as described in Section 2.1.1. However, there is no technical difficulty in the generalisation so that the proofs from [GPV2021] still apply, almost without modification.
We first present a utility result: computing all the numbers $D_{n,m,k}^{\mathcal{P}}$ up to a certain bound on $n$$m$, and $k$ can be done in polynomial time. This is of moderate interest in itself, but this is a requirement for our samplers, that compute theses values as a pre-processing step. Our algorithm is based on the so-called “recursive method” from [NW1978].
@@ 3.2 Recursive random sampling
A straightforward application of the recursive method from Nijenhuis and Wilf [NW1978] leads to Algorithm 1, which is presented here in a high level fashion.
@@
In [GPV2021, §3], we discuss how to implement Algorithm 1 efficiently. In particular we suggest a data-structure to represent DOAGs that allows for an efficient selection of the subset $I$ at line 8 and the function $f$ at line 9. In addition, implementation considerations are also given for the pick instruction at line 6, which is the core of the “recursive method”. As mentioned above, the numbers $D_{n,m,k}^{\mathcal{P}}$ either have to be pre-computed for Algorithm 1 to work, or must be lazily computed and memoised on the fly.
In practice, pre-computing all the necessary numbers to sample a uniform DOAG with $50$ vertices (without any constraint on $m$ and on the out-degree) using our library already takes about $8$ seconds on a standard laptop. This running time rapidly increases, which makes the cost generating large structures prohibitive. However, when limiting the number of edges and using a finite set $\mathcal{P}$, one can achieve much better results. For instance generating the four large bounded-degree DOAGs from Figure 3(b) takes about $11$ seconds on the same laptop, most of this time being spent in the pre-computation.
@@
Four random DOAGs, drawn using Algorithm 1 with $\mathcal{P}=\{0,1,2\}$$n=1250$$m=1300$ and $k=1$ are pictured in Figure 3(b). As a comparison, a truncated version of the Git history of the linux kernel is pictured on the left in Figure 3(a). Expectedly, the Git history looks more structured. This is because developers work on short-lived branches, consisting of chains of commits, generally starting from the main branch and merged back on main after the new feature (or bug fix, etc.) has been completed, reviewed and accepted.

### v4

# Asymptotic analysis and efficient random sampling of directed ordered acyclic graphs

@@ 1 Introduction
Directed Acyclic Graphs (DAGs for short) are directed graphs in which there is no directed path (sequence of incident edges) from a vertex to itself. They are an omnipresent data structure in various areas of computer science and mathematics. In concurrency theory for instance, scheduling problems usually define a partial order on a number of tasks, which is naturally encoded as DAG via its Hasse diagram [6, 4]: each task corresponds to a vertex in the graph and task dependencies are materialised by directed edges. Scheduling then corresponds to finding a good topological order on this graph. Natural question such as counting or sampling the schedulings of a program are studied in this context for the purpose of random testing [23]. DAGs also appear as the result of the compression of some tree-like structures such as XML documents [3]. In functional programming in particular, this happens at the memory layout level of persistent tree-like values, where the term “hash-consing” has been coined to refer to this compression [22]. Computer algebra systems also make use of this idea to store their symbolic expression [13]. These tree use cases leverage the fact that actual memory gains can be obtained by compacting trees, which has been quantified in [16] and has motivated the study of compacted trees in the recent years [12, 21]. Finally, complex histories, such as those used in version control systems (see Git for instance [17, p. 17]) or genealogy “trees” are DAGs as well.
@@ 1.1 Background on DAGs
The problem of counting DAGs has been solved in early 70’s by Robinson and Stanley using different approaches. In [38], Robinson exhibits a recursive decompositions of labelled DAGs leading to a recurrence satisfied by the numbers $A_{n,k}$ of DAGs with $n$ vertices including $k$ sources (vertices without any incoming edge). He later reformulates those results in terms of a new kind of generating functions, now called graphical generating functions in [36], and also obtains the asymptotic number of size $n$ DAGs. Around the same time, Stanley also used a generating function approach in [41] obtained the same results by deriving identities of the chromatic polynomial. Robinson also solves the unlabelled case starting from the same ideas but using Burnside’s lemma and cycle index sums to account for the symmetries of these objects. He provides a first solution in [38] and makes it more computationally tractable in [37]. In the 90’s, Gessel generalised those results, also using the graphical generating function framework in [20, 19] to take into account more parameters and count DAGs by vertices and edges, but also sinks and sources.
@@
From the point of view of uniform random generation, the recursive decomposition exhibited by Robinson in [36] is interesting as it is amenable to the recursive method pioneered by Nijenhuis and Wilf in [35]. This yields a polynomial time algorithm for sampling uniform DAGs with $n$ vertices. The analysis of this algorithm has been done in [31] but it had been acknowledged earlier in [34] although the article proposes an alternative solution. Both [31] and [34] also offer a Markov chain approach to the random sampling problem and an interesting discussion on the pros and cons of both approaches is given in [31]. Remote from the field of combinatorics, the random generation of DAGs is also an active topic in the area of applied statistics and Bayesian inference. In this context, DAGs encode a relevant structure in a collection of random variables and the problem of interest is to sample DAGs from a particular distribution related to those random variables. To this end, authors resort both to Monte Carlo Markov Chains approaches [32, 30] and methods similar to what is referred to as the recursive method in combinatorics [43]. An important point in [30] is better performance can be achieved by using a combination of both approaches, in particular by exploiting the combinatorial properties of DAGs. Notable is that sampling from the uniform distribution is tackled as a particular case in [43] and solved with the asymptotically optimal $O(n^{2})$ complexity at the expense of a $O(n^{3})$ pre-processing step.
@@
Another interesting question is that of controlling the number of edges in those random samplers. Indeed, sampling a uniform DAG with a prescribed number of vertices and edges cannot be achieved using the Markov chain approach as it constrains the chain too much, and the formulas of Gessel are not amenable to this either. In [31, § 7], the authors provide an interesting discussion on which kind of restrictions can be made on DAGs with the Markov chain approach. They address in particular the case of bounding the number of edges and highlight that the Monte Carlo Markov Chain approach fails when the desired number of edges is too low, thus advocating for having precise combinatorial enumerations.
@@ 1.2 Contributions
In the present paper, we propose to study an alternative model of DAGs, which we call Directed Ordered Acyclic Graphs (DOAGs), and which are enriched with additional structure on the edges. More precisely, a DOAG in an unlabelled DAG where (1) set of outgoing edges of each vertex is totally ordered and (2) the sources are totally ordered as well. This local ordering of the outgoing edges allows to capture more precisely the structure of existing mathematical objects. For instance, the compressed formulas and tree-like structures mentioned earlier (see [13, 22]) indeed present with an ordering as soon as the underlying tree representation is ordered. This is the case for most trees used in computer science (e.g. red-black trees, B-trees, etc.) and for all formulas involving non-commutative operators. The model we introduce thus allows for a more faithful modelling of a wide range of objects. We present here several results regarding DOAGs, as well as an extension of our method to classical labelled DAGs.
As a first step of our analysis, we describe a recursive decomposition scheme that allows us to study DOAGs using tools from enumerative combinatorics. This allows us to obtain a recurrence formula for counting them, as well as a polynomial-time uniform random sampler, based on the recursive method from [35], giving full control over their number of vertices and edges. Our decomposition is based on a “vertex-by-vertex” approach, that is we remove one vertex at a time and we are able to describe exactly what amount of information is necessary to reconstruct the graph. This differs from the approach of Robinson to study DAGs, where all the sources of a DAG are removed at once instead. Although this is a minor difference, our approach allows us to easily account for the number of edges of the graph, which is why our random sampler is able to target DOAGs with a specific number of edges. In terms of application, this means that we are able to efficiently sample DOAGs of low density. A second by-product of our approach is that it makes straightforward to bound the out-degree of each vertex, thus allowing to sample DOAGs of low degree as well.
In order to show the applicability of our method, we devise a similar decomposition scheme for counting labelled DAGs with any number of vertices, edges, and sources. This allows us to transfer our results on DOAGs in the context of labelled DAGs. Our new recurrence differs from the formula of Gessel [19] in that it does not resort to the inclusion-exclusion principle. Our approach allows us to obtain an efficient uniform random sampler of labelled DAGs with a prescribed number of vertices, edges, and sources. Here again, in addition to giving control over the number of edges of the produced objects, our approach can also be adapted to bound the out-degree of their vertices. To our knowledge, this is the first such sampler.
@@
In terms of applications, our random generation algorithms enable to experiment with the properties of the objects they model and with the average complexity of algorithms operating on them. A similar approach is for instance taken in [8] where samplers for a realistic class of Git graphs are developed in order to tackle the average complexity of a new algorithm introduced in [33, 7]. Random testing a also an important application of random sampling, especially as a building block for property-based testing, a now well-established framework pioneered Claessen and Hughes in [5].
This paper extends an earlier article [18] with new results on the asymptotics of DOAGs, with an optimal uniform random sampler for the case when the number of edges is not prescribed, and covers a larger class of DOAGs and DAGs by drooping a constraint on the number of sinks. For the sake of completeness, the most important results and ideas from [18] will be recalled in the present paper, but the reader will have to refer the earlier article to get the full proof and algorithmic details.
@@ 1.3 Outline of the paper
In Section 2, we start by introducing the class of Directed Ordered Acyclic Graphs and their recursive enumeration and describe a recursive decomposition scheme allowing to count them. In Section 3, we quickly go over earlier results regarding the random generation of DOAGs with a prescribed number of vertices, edges, and sources. The presentation given in this paper slightly generalises over the algorithm given in [18] but the ideas and proofs remain unchanged. Section 4 shows that our approach applies to labelled graphs as well and opens the way for further research regarding this class. We show that our method, when applied to labelled DAGs, yields a constructive counting formula for them, that is amenable to efficient uniform random generation with full control on the number of edges. Then, in Section 5, we present a bijection between DOAGs and class of integer matrices. This bijection is the key result of this paper as it allows to understand the structure of DOAGs in detail, and to obtain both asymptotic and algorithm results in the following sections. In Section 6, we present a first asymptotic result: we give an asymptotic equivalent of the number of DOAGs of size $n$ with any number of sources and edges. We also state some simple structural properties of those DOAGs. In light of the matrix encoding and these asymptotic results, we design an optimal uniform random sampler of DOAGs with a given number of vertices (but no constraint on the number of edges), that is described in Section 7.
@@ 2 Definition and recursive decomposition
In this section, we recall a model of directed acyclic graphs called “Directed Ordered Acyclic Graphs” (or DOAGs) that we introduced in [18]. It is similar to the classical model of unlabelled DAGs but where, in addition, we have a total order on the outgoing edges of each vertex. The presentation we opted for here slightly differs from that of [18] but essentially defines the same objects, the only difference being that we now allow several sinks for the sake of generality.
@@
  • for all $\in Vv\prec_{v}$ is a total order over the set of outgoing edges of $v$;
@@
We study this class as a whole, however, some sub-classes are also of special interest, in particular for the purpose of modelling compacted data structures. Tree structures representing real data, such as XML documents for instance [3], are rooted trees. When these trees are compacted, the presence of a root translates into a unique source in the resulting DOAG. Similarly, DOAGs with a single sink will arise naturally when compacting trees which bear a single type of leaves. In particular the model of compacted binary trees, which can also be seen as a class of cycle-free binary automata, has been shown have unusual combinatorial properties in [11, 12] and corresponds to a restriction of our model with only binary nodes (and one sink). For these reasons, we will also discuss how to approach the sub-classes of DOAGs with a single source and/or a single sink in this document.
@@
Figure 1: All DOAGs with respectly 11 vertex, 22 vertices, 33 vertices, and simultaneously 44 vertices and $3$ edges. All edges are implicitly oriented from top to bottom, the blue labels and arrows represent the sources and out-edges orderings (always from left to right).
@@ 2.1 Recursive decomposition
  • II is any subset of the internal vertices of $D^{\prime}$ (children of $v$ in $D$);
  • $s$ is any integer between 0 and the number of sources of $D^{\prime}$;
@@
[…]
where $p=s+i$ corresponds the out-degree of the smallest source. The term $\binom{n-k-p+i}{i}=\binom{n-k-s}{i}$ accounts for the choice of the set $I$ and the term $\binom{p}{i}i{!}$ accounts for the number of injective functions $f:I\to\llbracket 1;p\rrbracket$. The upper bound on $p$ in the sum is justified by the fact that the out-degree of any vertex can be at most the number of non-sources in the graph, that is (nk)(n-k).
The decomposition scheme presented here differs from the approach described by Robinson in [38] as it operates on only one source at a time. It is also reminiscent of the peeling processes used in map enumeration where maps are decomposed one face at a time, see for instance [27]. However, the absence of ordering amongst the incoming edges of each vertex in our setup renders those approaches inapplicable as is.
@@ 2.1.1 Special sub-classes based on out-degree constraints
Since $p=i+s$ is the out-degree of the removed source in the above summation, it is easy to adapt this sequence for counting DOAGs with constraints on the out-degree of the vertices. For instance, DOAGs with only one sink are obtained by ensuring that every vertex has out-degree at least one. In other words, let the summation start at $p=1$. Note that restricting DOAGs to have only one single sink or one single source ensures that they remain connected, however not all connected DOAGs are obtained this way. As another example, DOAGs with out-degree bounded by some constant $d$ are obtained by letting $p$ range from 0 to $\min(n-k,d)$.
@@
[…]
@@ 3 Earlier results on counting and recursive sampling
In this section we summarise our earlier results on the counting and random sampling problem for DOAGs when all three parameters (number of vertices, edges and sources) are fixed. The theorems are stated in a slightly more general setting here than in [18] so as to capture all variants of the model as described in Section 2.1.1. However, there is no technical difficulty in the generalisation so that the proofs from [18] still apply, almost without modification.
We first present a utility result: computing all the numbers $D_{n,m,k}^{\mathcal{P}}$ up to a certain bound on nnmm, and $k$ can be done in polynomial time. This is of moderate interest in itself, but this is a requirement for our samplers, that compute theses values as a pre-processing step. Our algorithm is based on the so-called “recursive method” from [35].
@@ 3.2 Recursive random sampling
A straightforward application of the recursive method from Nijenhuis and Wilf [35] leads to Algorithm 1, which is presented here in a high level fashion.
@@
In [18, §3], we discuss how to implement Algorithm 1 efficiently. In particular we suggest a data-structure to represent DOAGs that allows for an efficient selection of the subset $I$ at line 8 and the function $f$ at line 9. In addition, implementation considerations are also given for the pick instruction at line 6, which is the core of the “recursive method”. As mentioned above, the numbers $D_{n,m,k}^{\mathcal{P}}$ either have to be pre-computed for Algorithm 1 to work, or must be lazily computed and memoised on the fly.
In practice, pre-computing all the necessary numbers to sample a uniform DOAG with $n=50$ vertices (without any constraint on $m$ and on the out-degree) using our library already takes about $8$ seconds on a standard laptop. This running time rapidly increases, which makes the cost generating large structures prohibitive. However, when limiting the number of edges and using a finite set $\mathcal{P}$, one can achieve much better results. For instance generating the four large bounded-degree DOAGs from Figure 3(b) takes about $11$ seconds on the same laptop, most of this time being spent in the pre-computation.
@@
Four random DOAGs, drawn using Algorithm 1 with $\mathcal{P}=\{0,1,2\}$n=1250n=1250m=1300m=1300 and $k=1$ are pictured in Figure 3(b). As a comparison, a truncated version of the Git history of the linux kernel is pictured on the left in Figure 3(a). Expectedly, the Git history looks more structured. This is because developers work on short-lived branches, consisting of chains of commits, generally starting from the main branch and merged back on main after the new feature (or bug fix, etc.) has been completed, reviewed and accepted.